OpenStudy (anonymous):

whats the best way to learn the graphs of functions like sqrtx+5

OpenStudy (anonymous):

 I know that one is like this, I just don't know how to read it as such on my own

OpenStudy (anonymous):

You could get familiar with the base functions, and transformations

OpenStudy (abb0t):

memorize the graphs of the basic functions like \[\sqrt{x}, x^{2}, x ^{\frac{ 1 }{ 3 }}, \ln(x), \frac{ 1 }{ x },e ^{x}\] those are the basic graphs you should know. Anything that follows is just a shift. For instance \[x ^{2}+2\] is basically just a parabola shifted up to. So instead of starting at zero, the parabola begins at 2.

OpenStudy (anonymous):

ok, I'll try to do that. thanks